# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

### v2

@@ Abstract
Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent. Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative visual contexts from the original image, resulting in visual information loss and distorted distribution. Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative visual contexts. We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations), enabling selective suppression. By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation. Experiments on four benchmarks and six LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced. Code will be released upon acceptance.
Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: chongfeng@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
@@ 1 Introduction
Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors. The model used is LLaVA-1.5.
@@
Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al. (2023c); Wang et al. (2023), often generating text that contradicts the visual input. These hallucinations stem from various sources, with language priors Chen et al. (2025); Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b) being the most fundamental underlying cause. Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs. As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns. This often leads to a tendency to rely on memorized linguistic associations when responding to user queries. For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
@@
Methods for mitigating language priors in LVLMs can generally be classified into two categories. The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al. (2023); Wu et al. (2022); Chen et al. (2025), though such approaches often incur substantial training costs. In this work, we focus on plug-and-play, training-free methods. Many recent studies adopt contrastive decoding Li et al. (2023a), which constructs negative contexts to expose and suppress language priors. Some approaches directly perturb the input image Leng et al. (2024); Wang et al. (2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al. (2024b); Zhang et al. (2025). These methods typically compute the difference between the logit distributions obtained from the original and modified (i.e., negative) contexts to reduce the model’s reliance on language priors. Following the paradigm introduced in VCD Leng et al. (2024), these methods routinely modify the original image to generate negative visual contexts. However, this prevailing design choice raises a critical question: Is constructing negative contexts through perturbations of the original image truly the most effective strategy for inducing and mitigating language priors?
We argue that the answer is unequivocally negative. Constructing negative contexts based on the original image can lead to two main issues: (1) Visual information loss, where the negative visual context retains partial visual information, which may be degraded during contrastive decoding. (2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution. This analysis raises an important question: How can the above issues be avoided when constructing negative contexts?
In fact, simply taking another image to construct negative context is sufficient. This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image. As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis. Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for mitigating language priors in LVLMs. Instead of modifying the original image, CICD constructs the negative context using a different image that shares the same language priors. Our method avoids the two key issues present in previous approaches. Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context. Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias. The formal analysis of our method is presented in Sec. 3.2.
@@
Focusing on language priors, we further explore whether all such priors should be strictly eliminated. We argue that the answer is negative. Language priors can be divided into two categories: (1) essential priors, which are necessary for maintaining textual fluency, and (2) detrimental priors, which introduce information unrelated to the visual input. Clearly, only the detrimental priors should be removed. This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones? As shown in Fig.2, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content. Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.3. Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
We validate the effectiveness of CICD on four widely used benchmarks: CHAIR Rohrbach et al. (2018), POPE Li et al. (2023b), AMBER Wang et al. (2024a), and DetailCaps Ye et al. (2025). To assess the generalizability of our approach, we conduct experiments across six LVLMs from three different model families: InstructBLIP Dai et al. (2023), LLaVA-1.5 Liu et al. (2024c), LLaVA-Next Liu et al. (2024d), Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025). Experimental results show that CICD consistently reduces hallucinations by effectively mitigating language priors in LVLMs, outperforming state-of-the-art methods across all benchmarks and models. The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
@@
  • We conduct an in-depth analysis of existing contrastive decoding methods and identify their inherent limitations rooted in the reliance on the original image. We further verify the consistency of language priors across images and propose a stable approach for distinguishing essential priors from detrimental ones.
@@ 2 Related Work
Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Sec.A.
@@ 2.1 Hallucination in LVLMs
Hallucinations in LVLMs primarily concern cross-modal consistency Li et al. (2023c); Wang et al. (2023), which requires that the generated content remains faithful to the visual input. Among the various causes of hallucination, language priors Zhibo et al. (2023); Wu et al. (2022) are among the most fundamental. These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs. Several approaches Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs. Alternatively, recent training-free methods have explored contrastive decoding Li et al. (2023a); Chuang et al. (2024), typically constructing negative visual contexts based on the original image Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b); Zhang et al. (2025) in an attempt to elicit and suppress language priors.
In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image. Specifically, such modifications may lead to visual information loss, biased language priors, and distribution mismatch. We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD. Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
@@ 3 Preliminary
Figure 4: We input the target image and a distinct image into the LVLM separately. During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones. For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
@@ 3.1 Contrastive Decoding
Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$.
Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors that are independent of the visual input. As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence. This constitutes a fundamental cause of hallucinations. A commonly used approach to mitigate such priors is contrastive decoding Li et al. (2023a); Leng et al. (2024), which modifies the original image to construct a native visual context that elicits language priors. These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
@@ 4.1 Image Retrieval
Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships. These visual elements tend to exhibit minimal overlap across natural images. Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications. We provide further discussion in Sec.C.2. Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
@@ 4.2 Identifying Detrimental Priors
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.5. We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.2 and Fig.3. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
@@ 5 Experiment
Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length.
5.1 Experiment Settings
Benchmarks
We evaluate the performance of CICD across four widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks. These include two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
Evaluated LVLMs
We evaluate the generalizability of CICD across six LVLMs from three model families: InstructBLIP Dai et al. (2023), the LLaVA series (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)), and the Qwen series (Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025)). Unless stated otherwise, all models are implemented at the 7B scale.
Baselines
We compare CICD with four recent training-free methods that specifically target language priors: VCD Leng et al. (2024), ICD Wang et al. (2024c), IBD Zhu et al. (2024b), and DeFG Zhang et al. (2025). All of these methods construct negative visual contexts based on the original image. In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
5.2 Implementation Details
@@ 5.3 Main Results
Table 4: Results on DetailCaps. The maximum generation length is 512.
@@
Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
@@ 5.4 Effectiveness on More LVLMs
Further experiments will be discussed in Appendix B.
@@ 6 Conclusion
In this paper, we address the issue of language priors in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images. Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
@@ B.1 Benchmarks
assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO Lin et al. (2014) dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning.
@@
is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al. (2023), Gemini-1.5-Pro Team et al. (2024), and GPT-4o OpenAI et al. (2024) for evaluation. It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
@@
is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. POPE incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. Finally, it consists of 9,000 questions and 1,500 images.
@@
combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.

### v3

# Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models

@@ Abstract
Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent. Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative contexts from the original image, resulting in visual information loss and distorted distribution. Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative contexts. We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations). By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation. Experiments on 4 benchmarks and 6 LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced. Code will be released once accepted.
Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models
Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: fengchong@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
@@ 1 Introduction
Figure 1: A case of language prior. Adding the same prefix when captioning another image will lead to the same language priors. The model used is LLaVA-1.5.
@@
Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al. (2023c); Wang et al. (2023), often generating text that contradicts the visual input. These hallucinations stem from various sources, with language priors Chen et al. (2025); Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b) being the most fundamental underlying cause. Specifically, since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs. As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns. This often leads to a tendency to rely on memorized linguistic associations when responding to user queries. For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
Figure 2: The overlap ratio of image-related words between responses generated from negative contexts and those from regular inputs. Both VCD and ICD exhibit over 30% overlap, indicating that their negative contexts retain a considerable amount of essential visual information. In contrast, our method results in a nearly negligible overlap. The detailed analysis methodology can be found in Appendix. D.
Methods for mitigating language priors in LVLMs can generally be classified into two categories. The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al. (2023); Wu et al. (2022); Chen et al. (2025), though such approaches often incur substantial training costs. Another line of work explores plug-and-play, training-free methods, which are also the focus of our study. Among them, many recent studies adopt contrastive decoding Li et al. (2023a), which constructs negative contexts to reveal and suppress language priors. Some approaches directly perturb the input image Leng et al. (2024); Wang et al. (2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al. (2024b); Zhang et al. (2025). These methods typically compute the difference between the logit distributions obtained from the original and perturbed (i.e., negative) contexts to reduce the model’s reliance on language priors. Overall, following the paradigm introduced in VCD Leng et al. (2024), these methods routinely perturb the original inputs to generate negative contexts. However, this prevailing design choice raises a critical question: Is perturbing the original context to construct negative contexts truly the most effective strategy for inducing and mitigating language priors?
@@
We argue that the answer is unequivocally negative. Constructing negative contexts based on the original inputs can lead to two main issues: (1) Visual information loss, where the negative context retains partial visual information, which may be degraded during contrastive decoding. As shown in Fig. 2, compared to regular responses, those generated from negative contexts retain over 30% of the essential visual information, which is subsequently weakened during the contrastive decoding process. (2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution. This analysis raises an important question: How should negative contexts be constructed to avoid the above issue?
@@
In fact, simply taking a different image to construct negative context is sufficient. This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image. As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis. Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for suppressing language priors in LVLMs. Instead of perturbing the original inputs, CICD constructs the negative context using a different image that shares the same language priors. Our method avoids the two key issues present in previous approaches. Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context. As shown in Fig. 2, only a negligible amount (approximately 0.11%) of essential visual information is retained in our negative context. Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias. The formal analysis of our method is presented in Sec. 3.2.
Focusing on language priors, we further explore whether all such priors should be strictly eliminated. The answer is intuitively negative. Language priors can be divided into two categories: (1) Essential Priors, which are necessary for maintaining textual fluency, and (2) Detrimental Priors, which introduce information unrelated to the visual input. Clearly, only the detrimental priors should be removed. This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones? As shown in Fig.3, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content. Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.4. Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
Extensive experiments on 4 widely used benchmarks and 6 LVLMs across three model families demonstrate the effectiveness and generalizability of our method. By suppressing language priors, CICD consistently reduces hallucinations and achieves state-of-the-art performance across all benchmarks. The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
@@
  • We analyze existing contrastive decoding methods in depth and reveal their limitations stemming from reliance on the original image. To address this, we verify the consistency of language priors across images and propose a stable approach to distinguish essential priors from detrimental ones.
@@ 2 Related Work
Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Appendix.A.
@@ 2.1 Hallucination in LVLMs
Hallucinations in LVLMs primarily concern cross-modal consistency Li et al. (2023c); Wang et al. (2023), which requires that the generated content remains faithful to the visual input. Among the various causes of hallucination, language priors Zhibo et al. (2023); Wu et al. (2022) are the most fundamental. These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs. Several approaches Chen et al. (2025); Liu et al. (2024a); Gunjal et al. (2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs. Alternatively, recent training-free methods have explored contrastive decoding Li et al. (2023a); Chuang et al. (2024), typically constructing negative contexts based on the original image Leng et al. (2024); Wang et al. (2024c); Zhu et al. (2024b); Zhang et al. (2025) in an attempt to elicit and suppress language priors.
In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image. Specifically, such modifications may lead to visual information loss and distorted distribution. We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD. Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
@@ 3 Preliminary
Figure 5: The diagram of our method. We input the target image and a distinct image into the LVLM separately. During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones. For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
@@ 3.1 Contrastive Decoding
Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the model with parameters $\theta$.
Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors independent of the visual input. As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence. This constitutes a fundamental cause of hallucinations. A commonly used approach to mitigate such priors is contrastive decoding Li et al. (2023a); Leng et al. (2024), which modifies the original image to construct a native visual context that elicits language priors. These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
@@ 4.1 Image Retrieval
Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships. These visual elements tend to exhibit minimal overlap across natural images. Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications. We provide further discussion in Appendix.C.2. Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
@@ 4.2 Identifying Detrimental Priors
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.6. We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}_{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.3 and Fig.4. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors. Further discussion can be found in Appendix. C.3
@@ 5 Experiment
Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length. Cs and Ci represent CHAIRs and CHAIRi respectively.
@@
5.1 Experiment Settings
Benchmarks
We evaluate the performance of CICD across 4 widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks. These include two generative benchmarks (CHAIR Rohrbach et al. (2018) and DetailCaps Ye et al. (2025)), one discriminative benchmark (POPE Li et al. (2023b)), and one hybrid benchmark (AMBER Wang et al. (2024a)).
Evaluated LVLMs
We evaluate the generalizability of CICD across 6 LVLMs from three model families: InstructBLIP Dai et al. (2023), the LLaVA series (LLaVA-1.5 Liu et al. (2024c) and LLaVA-Next Liu et al. (2024d)), and the Qwen series (Qwen-VL-Chat Bai et al. (2023b), Qwen2-VL-Instruct Wang et al. (2024b), and Qwen2.5-VL-Instruct Bai et al. (2025)). Unless stated otherwise, all models are implemented at the 7B scale.
Baselines
We compare CICD with 4 recent training-free methods that specifically target language priors: VCD Leng et al. (2024), ICD Wang et al. (2024c), IBD Zhu et al. (2024b), and DeFG Zhang et al. (2025). All of these methods construct negative contexts based on the original image. In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
5.2 Implementation Details
@@ 5.3 Main Results
Table 4: Results on the COCO subset of DetailCaps. The maximum generation length is 512. Cs, Ci, and CAP represent CHAIRs, CHAIRi, and CAPTURE respectively.
@@
Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA).
@@ 5.4 Effectiveness on More LVLMs
Further experiments will be discussed in Sec. B.
@@ 6 Conclusion
In this paper, we mitigate the influence of language priors in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images. Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
@@ B.1 Benchmarks
Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
[…]
Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
Table 6: Results on the COCO subset of DetailCaps with more LVLMs.
@@
evaluates the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO Lin et al. (2014) dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning. This benchmark includes two metrics: CHAIRs and CHAIRi, defined as follows:
CHAIRs $=\frac{|\text{Hallucinated Objects}|}{|\text{All Objects}|},$ (7)
CHAIRi $=\frac{|\text{Hallucinated Sentences}|}{|\text{All Sentences}|}$
@@
is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al. (2023), Gemini-1.5-Pro Team et al. (2024), and GPT-4o OpenAI et al. (2024) for evaluation. It contains 4,870 images collected from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments. This benchmark primarily adopts the CAPTURE metric to evaluate caption quality. CAPTURE evaluates the alignment between generated and reference captions by calculating F1 scores using both hard and soft matching across entities ($F1_{obj}$), attributes ($F1_{attr}$), and relations ($F1_{rel}$). The final score is computed through a weighted aggregation of these components:
$CAPTURE=\frac{\alpha F1_{obj}+\beta F1_{attr}+\gamma F1_{rel}}{\alpha+\beta+\gamma}$ (8)
where $\alpha=5,\beta=5,\gamma=2$
@@
is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. It comprises three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. In total, it includes 9,000 questions and 1,500 images. Accuracy and F1 score are used as the primary evaluation metrics.
@@
combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition. AMBER contians mulitple metrics: CHAIR, Cover, Hal, Cog. It provides an annotated objects list $A_{obj}={obj_{1}^{A},obj_{2}^{A},\cdots,obj_{n}^{A}}$, and the generated objects are labeled as $R^{\prime}_{obj}$. Each metric is calculated as follows:
CHAIR $=1-\frac{\mathrm{len}(R^{\prime}_{obj}\cap A_{obj})}{\mathrm{len}(R^{\prime}_{obj})},$ (9)
Cover $=\frac{\mathrm{len}(R^{\prime}_{obj}\cap A_{obj})}{\mathrm{len}(A_{obj})},$
Hal =$=\frac{\{\text{CHAIR}>0\}}{\{\text{All Caps}\}},$
Cog $=\frac{\mathrm{len}(R^{\prime}_{obj}\cap H_{obj})}{\mathrm{len}(R^{\prime}_{obj})},$
where $H_{obj}$ denotes the set of hallucinated target objects generated by the LVLMs, and All Caps refers to all generated captions.